# CItruS [Uncaptioned image]: Chunked Instruction-aware State Eviction for Long Sequence Modeling

### v1

@@ Abstract
Long sequence modeling has gained broad interest as large language models (LLMs) continue to advance. Recent research has identified that a large portion of hidden states within the key-value caches of Transformer models can be discarded (also termed evicted) without affecting the perplexity performance in generating long sequences. However, we show that these methods, despite preserving perplexity performance, often drop information that is important for solving downstream tasks, a problem which we call information neglect. To address this issue, we introduce Chunked Instruction-aware State Eviction (CItruS), a novel modeling technique that integrates the attention preferences useful for a downstream task into the eviction process of hidden states. In addition, we design a method for chunked sequence processing to further improve efficiency. Our training-free method exhibits superior performance on long sequence comprehension and retrieval tasks over several strong baselines under the same memory budget, while preserving language modeling perplexity.
@@
Yu Bai1, Xiyuan Zou2,3thanks:   Equal contribution., Heyan Huang1 thanks:   Corresponding author., Sanxing Chen5, Marc-Antoine Rondeau2, Yang Gao1, and Jackie Chi Kit Cheung2,3,4 1Beijing Institute of Technology 2Mila – Quebec Artificial Intelligence Institute 3McGill University 4Canada CIFAR AI Chair 5Duke University
@@ 1 Introduction
Overall, our contributions are summarized as follows:
  1. 1.
    We define and demonstrate the information neglect problem in state-eviction methods.
  2. 2.
    We propose CItruS, a state eviction method designed for long sequence downstream tasks, which incorporates an instruction-aware cache for task-solving and a chunked state eviction process for efficient language modeling.
  3. 3.
    Experiments on long document reading comprehension, knowledge retrieval, and language modeling show that CItruS improves performance on downstream tasks involving long sequence by large margins while maintaining low language modeling perplexity.
@@ 2.1 Long Sequence Processing
Long sequence processing has long been a key research area in natural language processing (Tiezzi et al., 2024). Various approaches have been explored to address this challenge, including Longformer and State Space Models (Beltagy et al., 2020; Gu et al., 2022; Gu and Dao, 2023). Additionally, memory-augmented models use external memory to handle long sequences (Kuhn and De Mori, 1990; Wu et al., 2022; Bertsch et al., 2024; Lu et al., 2024), while recurrent-based transformers have been designed for long-sequence tasks (Dai et al., 2019; Li et al., 2023; Peng et al., 2023). More related work about long sequences is further provided in Appendix J.
@@
Our work is also similar to retrieval-augmented generation (RAG) methods (Gao et al., 2023; Zhao et al., 2024), which incorporates knowledge from external databases to enhance the generation. However, RAG research mainly focuses on the retrieval process in order to better leverage the documents that could support the response generation, whereas CItruS is a method that more generally focuses on performing various long sequence tasks. It could be a good option to be applied to the RAG process. In fact, our testing includes long-document question answering and retrieval as primary tasks.
@@ 5.1 Tasks
We compare the models using the following tasks. Detailed information about dataset statistics, prompts, and the divisions of document and instruction are provided in Appendices B and E.
@@
We use two tasks to test if the model could preserve the important information during the whole language modeling process: passkey retrieval [^3] https://huggingface.co/datasets/lvwerra/needle-llama3-16x524k (Mohtashami and Jaggi, 2023) and needle-in-a-haystack  [^4] https://github.com/gkamradt/LLMTest_NeedleInAHaystack tasks. The passkey retrieval task tests if the model can retrieve a single passkey (e.g., a five-digit number) inserted in a synthetic long document made up by repetitive simple sentences. We conduct this task on the documents with lengths up to 1 million tokens. The needle-in-a-haystack task replaces the passkey with a more general text fact and inserts them in real long documents. An example of the fact and the information of the documents can be found in Appendix F. The maximum length of documents for needle-in-a-haystack is set to 12,000. We use accuracy in the passkey retrieval task and the ROUGE metric (Lin, 2004) for the needle-in-a-haystack task to award partial correctness.
@@
We report perplexity scores on long-range language modeling to estimate how well our models maintain fluency in generation (Xiao et al., 2023). We used PG19 (Rae et al., 2019) dataset.
@@ 5.3 Hyperparameters
We applied the position shift mechanism leveraged by Xiao et al. (2023), which always use the same positional embeddings for the caches containing different hidden states, to make the models process long documents better. We also apply this technique to all the baselines to enhance their ability of processing long sequences. We use the Llama 2 Chat model (Touvron et al., 2023) with 7 billion and 13 billion parameters and the 7 billion parameter Mistral Instruct model (Jiang et al., 2023) as the backbone models. $k$ is set as $768$ and $l_{s}$ is set as $256$, resulting a cache size of $1{,}024$ during modeling the document. This setting is also applied to all the baseline models. We apply 8 bit quantization on the 13 billion parameter model. Results are inferred on one A100 80G GPU. All the hyperparameters are selected using the validation sets.
@@ 6.3 Long-range language modeling
We compare our model with the long-range language modeling model, Streaming LLM. Specifically, we evaluate the standard CSE as well as the shared cache version of our proposed CItruS. For CItruS with a shared cache, we randomly sample 10 different instructions including different questions from Qasper and HotpotQA dataset. We show the results using one instruction here and append the rest of the results in the Appendix G. Results in Figure 7 show that our standard CSE could maintain the perplexity when processing long sequences as low as the Streaming LLM. Meanwhile, although showing a slight increase in perplexity with the Llama 2 7B Chat model, CSE with a shared cache achieves consistent perplexity results without exploding as described by Xiao et al. (2023). This shows that introducing the instruction text as the query to evict hidden states would not affect the text perplexity of the large language models. A more detailed discussion about the roles of the standard cache and the instruction-aware cache in our model is provided in Appendix H.
@@ 6.4 Analysis
In this section, we provide analyses on the hyper-parameters of our model, the effect of chunk size, and the position bias in the knowledge retrieval tasks. We also provide an analysis on the effect of the initial tokens in Appendix I. We report the averaged results in this section since all the models perform similarly in the analyses across all the datasets. The full results are shown in Appendix C.
@@ Appendix D Information Neglect of the Sliding Window Methods
Appendix E Prompt for Each Task
@@
Appendix F Setup for the Needle-in-a-Haystack Task
@@
Appendix G Results of Perplexity with Other Instructions
@@ Appendix H Discussion
Appendix I Analysis on initial tokens
@@
Appendix J More Related Work
J.1 Long Sequence Processing
Long sequence language modeling have attracted more and more research interests in recent years (Tiezzi et al., 2024). Various long document processing tasks are proposed to evaluate the long sequence modeling of language models (Zhao et al., 2021; Luo et al., 2021; Bai et al., 2023). Longformer, leveraging sparse self-attention pattern, save the memory cost to make the model process long document (Beltagy et al., 2020). Memorizing transformer uses a external memory to save the information during the long sequence modeling process (Wu et al., 2022). Mistral applied Pre-fill and chunking sliding window methods to model longer sequences  (Jiang et al., 2023). State space models and their variations are also popular recently (Gu et al., 2022; Gu and Dao, 2023; Wang et al., 2022). Unlimitedformer wraps pretrained encoder-decoder transformer, and offloads the cross-attention computation to a single k-nearest-neighbor index, while the returned $k$NN distances are the attention dot-product scores (Bertsch et al., 2024).  Nawrot et al. (2024) propose to compress the key-value cache to make the model process longer sequences. Xiong et al. (2023) conduct continual pretraining from Llama 2 (Touvron et al., 2023) with longer training sequences and on a dataset where long texts are upsampled. Rotary Position Embedding and the positional interpolation based on it are also used enable the model process longer sequences (Su et al., 2024; Chen et al., 2023). ReadAgent are proposed by using a large language model agent to process long sequences (Lee et al., 2024). LongHeads enhances the long-context processing of large language models by allowing multi-head attention to attend to selected important context chunks within the trained length (Lu et al., 2024). Infini-Transformer leverage a compressive memory between different context segment to achieve modeling long range text (Munkhdalai et al., 2024). Hwang et al. (2024) propose TransformerFAM, a novel architecture with a feedback loop for attending to latent representations, enables Transformers to process indefinitely long sequences without additional weights. Zhang et al. (2024a) leverage plug-and-play positional encoding to make the model better collect the information in the middle of the document.
@@
J.2 State Eviction for Large Language Models

### v2

@@ Abstract
Long sequence modeling has gained broad interest as large language models (LLMs) continue to advance. Recent research has identified that a large portion of hidden states within the key-value caches of Transformer models can be discarded (also termed evicted) without affecting the perplexity performance in generating long sequences. However, we show that these methods, despite preserving perplexity performance, often drop information that is important for solving downstream tasks, a problem which we call information neglect. To address this issue, we introduce Chunked Instruction-aware State Eviction (CItruS), a novel modeling technique that integrates the attention preferences useful for a downstream task into the eviction process of hidden states. In addition, we design a method for chunked sequence processing to further improve efficiency. Our training-free method exhibits superior performance on long sequence comprehension and retrieval tasks over several strong baselines under the same memory budget, while preserving language modeling perplexity. The code and data have been released at https://github.com/ybai-nlp/CItruS.
@@
Yu Bai1,2, Xiyuan Zou3,4thanks:   Equal contribution. Work done during the internship at Mila – Quebec Artificial Intelligence Institute., Heyan Huang1,2 thanks:   Corresponding author., Sanxing Chen6, Marc-Antoine Rondeau3, Yang Gao1, and Jackie Chi Kit Cheung3,4,5 1Beijing Institute of Technology, Beijing, China 2Southeast Academy of Information Technology, Fujian, China 3Mila – Quebec Artificial Intelligence Institute 4McGill University    5Canada CIFAR AI Chair    6Duke University yubai@bit.edu.cn, xiyuan.zou@mail.mcgill.ca
@@ 1 Introduction
Overall, our contributions are summarized as follows: 1) We define and demonstrate the information neglect problem in state-eviction methods. 2) We propose CItruS, a state eviction method designed for long sequence downstream tasks, which incorporates an instruction-aware cache for task-solving and a chunked state eviction process for efficient language modeling. 3) Experiments on long document reading comprehension, knowledge retrieval, and language modeling show that CItruS improves performance on downstream tasks involving long sequence by large margins while maintaining low language modeling perplexity.
@@ 2.1 Long Sequence Processing
Long sequence processing has long been a key research area in natural language processing (Tiezzi et al., 2024). Various approaches have been explored to address this challenge, including Longformer and State Space Models (Beltagy et al., 2020; Gu et al., 2022; Gu and Dao, 2023). Additionally, memory-augmented models use external memory to handle long sequences (Kuhn and De Mori, 1990; Wu et al., 2022; Bertsch et al., 2024; Lu et al., 2024), while recurrent-based transformers have been designed for long-sequence tasks (Dai et al., 2019; Li et al., 2023; Peng et al., 2023). More related work about long sequences is further provided in Appendix M.
@@
Our work is also similar to retrieval-augmented generation (RAG) methods (Gao et al., 2023; Zhao et al., 2024), which incorporates knowledge from external databases to enhance the generation. However, RAG research mainly focuses on the retrieval process in order to better leverage the documents that could support the response generation, whereas CItruS is a method that more generally focuses on performing various long sequence tasks. It could be a good option to be applied to the RAG process. In fact, our testing includes long-document question answering and retrieval as primary tasks. We further discuss the difference between our method and RAG in Appendix M.
@@ 3 The Information Neglect Problem
Additionally, we conduct another set of experiments that demonstrate the performance degradation of the standard state eviction models compared to the standard models that use the full text as input. Results supporting the presence of information neglect problem are presented in Appendix E.
@@ 5.1 Tasks
We compare the models using the following tasks. Detailed information about dataset statistics, prompts, and the divisions of document and instruction are provided in Appendices B and F.
@@
We use two tasks to test if the model could preserve the important information during the whole language modeling process: passkey retrieval [^3] https://huggingface.co/datasets/lvwerra/needle-llama3-16x524k (Mohtashami and Jaggi, 2023) and needle-in-a-haystack  [^4] https://github.com/gkamradt/LLMTest_NeedleInAHaystack tasks. The passkey retrieval task tests if the model can retrieve a single passkey (e.g., a five-digit number) inserted in a synthetic long document made up by repetitive simple sentences. We conduct this task on the documents with lengths up to 1 million tokens. The needle-in-a-haystack task replaces the passkey with a more general text fact and inserts them in real long documents. An example of the fact and the information of the documents can be found in Appendix G. The maximum length of documents for needle-in-a-haystack is set to 12,000. We use accuracy in the passkey retrieval task and the ROUGE metric (Lin, 2004) for the needle-in-a-haystack task to award partial correctness. Additional experiments using BABILong (Kuratov et al., 2024), a dataset design for the long-context needle-in-a-haystack task, are also conducted in Appendix I.
@@
We report perplexity scores on long-range language modeling to estimate how well our models maintain fluency in generation (Xiao et al., 2023). We used PG19 (Rae et al., 2019) dataset for this task.
@@ 5.3 Hyperparameters
We applied the position shift mechanism leveraged by Xiao et al. (2023), which always use the same positional embeddings for the caches containing different hidden states, to make the models process long documents better. We also apply this technique to all the baselines to enhance their ability of processing long sequences. We use the Llama 2 Chat model (Touvron et al., 2023) with 7 billion and 13 billion parameters and the 7 billion parameter Mistral Instruct model (Jiang et al., 2023) as the backbone models. Additionally, we include experiments using Llama 3 8B Instruct model, shown in Appendix H. $k$ is set as $768$ and $l_{s}$ is set as $256$, resulting a cache size of $1{,}024$ during modeling the document. This setting is also applied to all the baseline models. We apply 8 bit quantization on the 13 billion parameter model. Results are inferred on one A100 80G GPU. All the hyperparameters are selected using the validation sets.
@@ 6.3 Long-range language modeling
We compare our model with the long-range language modeling model, Streaming LLM. Specifically, we evaluate the standard CSE as well as the shared cache version of our proposed CItruS. For CItruS with a shared cache, we randomly sample 10 different instructions including different questions from Qasper and HotpotQA dataset. We show the results using one instruction here and append the rest of the results in the Appendix J. Results in Figure 7 show that our standard CSE could maintain the perplexity when processing long sequences as low as the Streaming LLM. Meanwhile, although showing a slight increase in perplexity with the Llama 2 7B Chat model, CSE with a shared cache achieves consistent perplexity results without exploding as described by Xiao et al. (2023). This shows that introducing the instruction text as the query to evict hidden states would not affect the text perplexity of the large language models. A more detailed discussion about the roles of the standard cache and the instruction-aware cache in our model is provided in Appendix K.
@@ 6.4 Analysis
In this section, we provide analyses on the hyper-parameters of our model, the effect of chunk size, and the position bias in the knowledge retrieval tasks. We also provide an analysis on the effect of the initial tokens in Appendix L. We report the averaged results in this section since all the models perform similarly in the analyses across all the datasets. The full results are shown in Appendix C.
@@ 7 Conclusion
Acknowledgements
The authors thank all the reviewers for their suggestions and comments. This work is supported by National Natural Science Foundation of China (No. U21B2009) and McGill Science Undergraduate Research Award (SURA). Jackie Chi Kit Cheung is supported by Canada CIFAR AI Chair program. The authors also acknowledge the material support of NVIDIA in the form of computational resources.
@@ Appendix D Information Neglect of the Sliding Window Methods
Appendix E Supplemental Experiments for Information Neglect Problem
We discussed the issue of information neglect in Section 3. In this section, we present a straightforward experiment to further demonstrate the existence of this problem. Specifically, we compare the performance of models that read the full context of the document with those employing state eviction techniques. This experiment utilizes the Llama 3 8B Instruct model across the six reading comprehension datasets mentioned in our paper. As most long documents exceed the model’s processing capacity, we limited our tests to examples with fewer than 4096 tokens. Additionally, we applied 8-bit quantization for efficiency. Alongside the previously discussed state eviction models, we also include our proposed CItruS model. We set $k=256,l_{s}=256$ for these experiments to simulate scenarios with small caches and long documents. The results are shown in Table 11:
Cache Type Avg. Rank Avg. Results Qasper Multifieldqa_En Hotpot QA TREC TriviaQA SamSum
Streaming LLM 2.83 33.94 23.58 16.25 38.27 44.29 46.39 34.85
TOVA 3.00 36.07 15.50 17.60 44.84 47.14 57.10 34.26
RoCo 2.50 34.00 21.83 25.47 28.33 42.86 51.93 33.58
H2O 4.33 37.96 16.39 25.69 40.20 54.29 62.17 29.03
Standard CSE 3.33 35.84 15.87 18.48 39.02 44.29 60.25 37.12
   $+$ Individual Cache 7.00 49.12 23.17 46.43 50.95 47.14 88.89 38.12
   $+$ Shared Cache 7.67 49.82 26.73 47.61 49.90 48.57 88.89 37.20
H2O $+$ Shared Cache 5.83 45.97 18.63 39.44 48.06 48.57 85.44 35.66
Full Text 7.67 49.87 23.97 52.29 46.56 54.29 83.45 38.65
Table 11: The average results and the average reversed ranks of the Llama 3 8B Instruct model on six different long sequence tasks, where Avg. Rank represents the averaged reversed rank and Avg. Results represents the averaged results.
Results show: 1) There is a large gap between the performance of the previous cache eviction methods and the model that could “read” the full text. 2) We would like to point out that this is not the ideal case for our proposed CItruS, which is designed for processing long sequences beyond the capacity of LLMs. However, even with the short context, the proposed method approaches the performance of full-context models better than the baseline models. Notably, in the TriviaQA and Qasper datasets, CItruS outperforms the models with the full text. We hypothesize that it is because some noisy information is eliminated during the eviction process.
Appendix F Prompt for Each Task
@@
Appendix G Setup for the Needle-in-a-Haystack Task
@@
Cache Type 0-4k 4-8k 8k+
Streaming LLM 2.17 2.33 1.83
TOVA 2.00 2.33 3.83
Roco 3.67 4.67 5.33
H2O 3.83 2.67 3.00
H2O $+$ Shared Cache 6.67 6.33 4.83
Standard CSE 3.33 3.00 3.33
   $+$ Individual Cache 6.67 7.00 7.17
   $+$ Shared Cache 7.17 7.33 6.33
Table 14: The averaged reversed rank results among all the 8 models on six different reading comprehension tasks with Llama 3 8B Instruct, where 8 is the highest score and 1 is the lowest score. Results are presented by grouping text with lengths of 0-4k, 4k-8k, and 8k$+$. Best results are bolded.
Cache Type 0-4k 4-8k 8k+
Streaming LLM 39.75 38.82 34.20
TOVA 41.09 39.88 38.65
Roco 45.17 45.10 42.41
H2O 45.45 42.24 39.13
H2O $+$ Shared Cache 50.85 51.86 48.24
Standard CSE 44.82 41.48 37.17
   $+$ Individual Cache 50.66 52.10 53.03
   $+$ Shared Cache 51.17 53.20 51.66
Table 15: The averaged results among all the 8 models on six different reading comprehension tasks with Llama 3 8B Instruct, where 8 is the highest score and 1 is the lowest score. Results are presented by grouping text with lengths of 0-4k, 4k-8k, and 8k$+$. Best results are bolded.
Appendix H Experimental Results with Llama 3
We test the six reading comprehension tasks used in our paper with the newly released Llama 3 8B Instruct model. The results shown in Table 14 and Table 15 demonstrates that our proposed model continues to perform consistently better than other state eviction models.
Appendix I Experimental Results with BABILong Dataset
We conduct supplementary experiments with BABILong (Kuratov et al., 2024), a newly proposed dataset which contains long sequence needle-in-the-haystack tasks involving multiple supporting facts and requires the model to generate answers using multi-hop reasoning and temporal dynamics.
We test our models and the baselines on the qa1, qa2, and qa3 subsets of BABILong with a maximum length of 128k tokens. All results were obtained using the Llama 3 8B Instruct model. The results are shown in Table 16, Table 17, and Table 18, where the “0k”, “1k”, and “64k” represent the context length of the subset.
Results show that our method performs better on these tasks, especially when the context length is longer. However, we want to point out that it is not guaranteed that our method could enhance the reasoning ability of LLMs. We are only claiming that our method can better help the state eviction methods retain more relevant information for downstream tasks when processing long sequences. The reasoning abilities depend on the model and how it leverages the information in the retained hidden states, which is fundamentally influenced by the pretraining process.
Model qa1_0k qa1_1k qa1_2k qa1_4k qa1_8k qa1_16k qa1_32k qa1_64k qa1_128k Avg_qa1
Streaming LLM 0.98 0.83 0.74 0.43 0.25 0.16 0.08 0.03 0.03 0.39
TOVA 0.98 0.83 0.68 0.48 0.41 0.29 0.15 0.04 0.02 0.43
Roco 0.98 0.83 0.60 0.51 0.31 0.13 0.04 0.05 0.01 0.38
H2O 0.98 0.83 0.31 0.20 0.13 0.05 0.01 0.01 0.01 0.28
H2O $+$ Shared Cache 0.98 0.90 0.87 0.75 0.62 0.51 0.28 0.16 0.10 0.57
Standard CSE 0.98 0.90 0.69 0.53 0.39 0.30 0.21 0.09 0.03 0.46
   $+$ Individual Cache 0.98 0.89 0.82 0.73 0.63 0.56 0.50 0.31 0.26 0.63
   $+$ Shared Cache 0.98 0.89 0.84 0.79 0.75 0.69 0.59 0.66 0.43 0.74
Table 16: The results on BABILong qa1 subset. Results are evaluated on test examples with different lengths. Best results are bolded.
Model qa2_0k qa2_1k qa2_2k qa2_4k qa2_8k qa2_16k qa2_32k qa2_64k qa2_128k Avg_qa2
Streaming LLM 0.14 0.14 0.39 0.23 0.08 0.01 0.01 0.00 0.00 0.11
TOVA 0.14 0.12 0.21 0.27 0.18 0.09 0.07 0.04 0.02 0.13
Roco 0.14 0.10 0.02 0.29 0.11 0.05 0.03 0.01 0.01 0.08
H2O 0.14 0.10 0.00 0.02 0.00 0.02 0.00 0.01 0.00 0.03
H2O $+$ Shared Cache 0.14 0.12 0.11 0.05 0.01 0.00 0.00 0.00 0.00 0.05
Standard CSE 0.14 0.13 0.12 0.27 0.23 0.10 0.05 0.04 0.02 0.12
   $+$ Individual Cache 0.14 0.12 0.13 0.15 0.26 0.24 0.28 0.26 0.24 0.20
   $+$ Shared Cache 0.14 0.12 0.19 0.12 0.11 0.04 0.12 0.14 0.25 0.14
Table 17: The results on BABILong qa2 subset. Results are evaluated on test examples with different lengths. Best results are bolded.
Model qa3_0k qa3_1k qa3_2k qa3_4k qa3_8k qa3_16k qa3_32k qa3_64k qa3_128k Avg_qa3
Streaming LLM 0.25 0.26 0.27 0.32 0.31 0.10 0.07 0.02 0.00 0.18
TOVA 0.25 0.24 0.21 0.40 0.31 0.19 0.10 0.06 0.01 0.20
Roco 0.25 0.26 0.02 0.19 0.23 0.12 0.07 0.04 0.01 0.13
H2O 0.24 0.23 0.01 0.01 0.00 0.00 0.00 0.01 0.00 0.06
H2O $+$ Shared Cache 0.24 0.23 0.22 0.07 0.04 0.02 0.02 0.01 0.01 0.10
Standard CSE 0.25 0.23 0.19 0.32 0.28 0.16 0.11 0.07 0.04 0.18
   $+$ Individual Cache 0.24 0.22 0.25 0.20 0.25 0.25 0.25 0.25 0.21 0.24
   $+$ Shared Cache 0.24 0.22 0.26 0.19 0.19 0.22 0.32 0.33 0.34 0.26
Table 18: The results on BABILong qa3 subset. Results are evaluated on test examples with different lengths. Best results are bolded.
Appendix J Results of Perplexity with Other Instructions
@@ Appendix I Experimental Results with BABILong Dataset
Appendix L Analysis on initial tokens
@@
Appendix M More Related Work
M.1 Long Sequence Processing
Long sequence language modeling have attracted more and more research interests in recent years (Tiezzi et al., 2024), as large language models continue to advance (Li et al., 2024a). Various long document processing tasks are proposed to evaluate the long sequence modeling of language models (Zhao et al., 2021; Luo et al., 2021; Bai et al., 2023). Longformer, leveraging sparse self-attention pattern, save the memory cost to make the model process long document (Beltagy et al., 2020). Memorizing transformer uses a external memory to save the information during the long sequence modeling process (Wu et al., 2022). Mistral applied Pre-fill and chunking sliding window methods to model longer sequences  (Jiang et al., 2023). State space models and their variations are also popular recently (Gu et al., 2022; Gu and Dao, 2023; Wang et al., 2022). Unlimitedformer wraps pretrained encoder-decoder transformer, and offloads the cross-attention computation to a single k-nearest-neighbor index, while the returned $k$NN distances are the attention dot-product scores (Bertsch et al., 2024).  Nawrot et al. (2024) propose to compress the key-value cache to make the model process longer sequences. Xiong et al. (2023) conduct continual pretraining from Llama 2 (Touvron et al., 2023) with longer training sequences and on a dataset where long texts are upsampled. Rotary Position Embedding and the positional interpolation based on it are also used enable the model process longer sequences (Su et al., 2024; Chen et al., 2023). Text summarization has also been known by its relation with long sequence processing area (Du and Gao, 2023; Gao et al., 2024; Li et al., 2024b). ReadAgent are proposed by using a large language model agent to process long sequences (Lee et al., 2024). LongHeads enhances the long-context processing of large language models by allowing multi-head attention to attend to selected important context chunks within the trained length (Lu et al., 2024). Infini-Transformer leverage a compressive memory between different context segment to achieve modeling long range text (Munkhdalai et al., 2024). Hwang et al. (2024) propose TransformerFAM, a novel architecture with a feedback loop for attending to latent representations, enables Transformers to process indefinitely long sequences without additional weights. Zhang et al. (2024a) leverage plug-and-play positional encoding to make the model better collect the information in the middle of the document.
@@
Retrieval-augmented generation techniques also share similar aspects with our methods. RAG techniques usually involve two steps: first, retrieving relevant information (usually a document) from a large database, and second, concatenating the document and the user query to enhance the performance of generating the response text. The similarity between our method and RAG methods mainly lies in the fact that our method can be applied to long document question-answering tasks, which is the typical form of the final step of the RAG methods. In this sense, our method is orthogonal to them, as it aims to improve the LLMs themselves and can handle documents that exceed the length limitations of LLMs in the RAG process. Hence, it is not appropriate to directly compare our methods to RAG techniques.
M.2 State Eviction for Large Language Models